Reece:

Hi, can someone please help me answer this question? Determine the exact value of each expression: cos (sin inverse (-square root of 3/2))

supie:

@darkknight

Reece:

\[\cos (\sin^{-1}( -\sqrt{3}/2))\]

Reece:

this is the problem there are no diagrams

supie:

@jhonyy9 ?

jhonyy9:

ok so we need get the invers of sinus function what is arcsin y = sin(x) => x = sin(y) => y = arcsin(x) \[\cos(\sin^{-1}(-\frac{ \sqrt{3} }{ 2 }))\] We want the value of arcsin(−sqrt3/2) in terms of π. Let arcsin(−sqrt3/2)=y ⇒siny=−sqrt3/2⇒−siny=sqrt3/2 sin(−y)=−siny⇒sin(−y)=sqrt3/2=sin(π/3) ⇒−y=π/3⇒y=−π/3 ⇒arcsin(−sqrt3/2)=−π/3 sin(π/3)=sqrt3/2 sin(π+π/3)=−sin(π/3)=−sqrt3/2 ⇒sin(π+π/3)=sin(4π/3)=−sqrt3/2 ⇒arcsin(−sqrt3/2)=4π/3 ⇒arcsin(−sqrt3/2)=−π/3 or 4π/3 and now we have again \[\cos (\frac{ -\pi }{ 3 }) = - \cos(\frac{ \pi }{ 3 }) = - \frac{ 1 }{ 2 }\]

jhonyy9:

@supie

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok so we need get the invers of sinus function what is arcsin y = sin(x) => x = sin(y) => y = arcsin(x) \[\cos(\sin^{-1}(-\frac{ \sqrt{3} }{ 2 }))\] We want the value of arcsin(−sqrt3/2) in terms of π. Let arcsin(−sqrt3/2)=y ⇒siny=−sqrt3/2⇒−siny=sqrt3/2 sin(−y)=−siny⇒sin(−y)=sqrt3/2=sin(π/3) ⇒−y=π/3⇒y=−π/3 ⇒arcsin(−sqrt3/2)=−π/3 sin(π/3)=sqrt3/2 sin(π+π/3)=−sin(π/3)=−sqrt3/2 ⇒sin(π+π/3)=sin(4π/3)=−sqrt3/2 ⇒arcsin(−sqrt3/2)=4π/3 ⇒arcsin(−sqrt3/2)=−π/3 or 4π/3 and now we have again \[\cos (\frac{ -\pi }{ 3 }) = - \cos(\frac{ \pi }{ 3 }) = - \frac{ 1 }{ 2 }\] \(\color{#0cbb34}{\text{End of Quote}}\) @XioGonz

jhonyy9:

@Vocaloid pls do you agree this ? ty.